OpenStudy (anonymous):

solve this system of equation using the addition method? x + y = 6 -x + 3y = -2

OpenStudy (anonymous):

(x - x) + (y + 3y) = (6 - 2) 4y = 4 y = 4 Place the value on any equation: x + y = 6 x + 4 = 6 x = -2

OpenStudy (anonymous):

y=1 x= 5

OpenStudy (anonymous):

you are using the termination method in thisprob. I added the equations the result was 4y=4 so y equals 1